OpenStudy (anonymous):

In triangle ABC,angle C=120 degree, then the value of cos^2 A + cos^2 B -cosA.cosB= A 3/4 B 3/2 C 1/2 D 1/4

ganeshie8 (ganeshie8):

\[A+B+120= 180 \implies A+B = 60 \]

ganeshie8 (ganeshie8):

\[\cos^2A + \cos^2B - \cos A\cos B = \left(\cos A -\cos B\right)^2 + \cos A \cos B\]

ganeshie8 (ganeshie8):

use cosC - cosD, and 2cosAcosB identities ^

OpenStudy (anonymous):

Yup I got it.... I had done that initially too but was not getting the answer... Thanxx a lot.i have a two more questions... Could help me out with them too?

ganeshie8 (ganeshie8):

I'l try, ask..

OpenStudy (anonymous):

The number of real solutions of the equation sin(e^x)=2^x+ 2^(-x) is- A 1 B 0 C 2 D none

ganeshie8 (ganeshie8):

hint : \(-1 \le \sin \theta \le 1\)

ganeshie8 (ganeshie8):

\(\large \sin(e^x) = 2^x + 2^{-x}\) \(\large \sin(e^x) = 2^x +\dfrac{1}{2^x}\) \(\large 2^x\sin(e^x) = 2^{2x} + 1\) \(\large 2^x[4-\sin(e^x) ] = -1\)

OpenStudy (anonymous):

It should be zero right? Because all the values of the result are greater than zero?

ganeshie8 (ganeshie8):

Clearly the left side is always positive, but the right side is negative - so no solutions.

OpenStudy (anonymous):

The left hs can be positive or negative right coz the power is on e whereas the rhs can only be positive because anything to the power anything cannot be lesser than zero

ganeshie8 (ganeshie8):

wait there is a mistake, \(2^{2x} \ne 4*2^x\)

OpenStudy (vishweshshrimali5):

There is another way to deal with your question.

OpenStudy (anonymous):

Yup. Basically, sina cannot be greater than or lesser than 1 but the hrs sums up to be always greater than 1 right?

OpenStudy (anonymous):

So number of solutions are zero

OpenStudy (vishweshshrimali5):

Yep sorry I have applied opposite inequalities sign in my statements. Here is the correct one.

OpenStudy (vishweshshrimali5):

\(2^x + \cfrac{1}{2^x} \ge 2\) or \(2^x + \cfrac{1}{2^x} \le -2 \) But, since \(-1 \le \sin (e^x) \le 1\) So there are no solutions for given equation.

OpenStudy (vishweshshrimali5):

Did you get this one ?

OpenStudy (anonymous):

Thanks a lot @vishweshshrimali5 and @ganeshie8

OpenStudy (vishweshshrimali5):

No problem.

OpenStudy (anonymous):

By the way can you help me with another question?

OpenStudy (vishweshshrimali5):

ya sure

OpenStudy (anonymous):

If x1 and x2 are two values lying in [0,360] degrees range,for which Tanx=p, tan((x1)/2).tan((x2)/2)= what?

OpenStudy (vishweshshrimali5):

Sorry is that tan(x) = pi

OpenStudy (anonymous):

No it's just a constant. In the question it was given lambda but I did not know how to represent it so p

OpenStudy (vishweshshrimali5):

No problem.

OpenStudy (vishweshshrimali5):

Also, in the given domain : \[\tan x _{1} = \tan (180 + x_{1})\] So \[x_{2} = 180 + x_{1}\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (vishweshshrimali5):

There was no need of that previous formula. You can use only the above result.

OpenStudy (vishweshshrimali5):

Which result are u talking about ?

OpenStudy (anonymous):

I know the previous formula . I do not know how you arrived to the second conclusion

OpenStudy (vishweshshrimali5):

Ohh: See we know that : \[\tan (\theta) = \tan(\pi + \theta) = \tan (2\pi + \theta) = ...\]

OpenStudy (vishweshshrimali5):

But because domain is only \([0,2\pi]\) so only possibility that two angles \(x_{1}\) and \(x_{2}\) to exist such that \[\tan (x_{1}) = \tan (x_{2}) = \lambda\] is \[x_{2} = x_{1} + \pi\]

OpenStudy (anonymous):

Ok

OpenStudy (vishweshshrimali5):

Now you have to find out the value of : \[\tan(\cfrac{x_{1}}{2}) \tan(\cfrac{x_{2}}{2})\] Now put \(x_{2} = \pi + x_{1}\) and use this result : \[\tan(x+\cfrac{\pi}{2}) = -\cot(x) = -\cfrac{1}{\tan(x)}\]

OpenStudy (anonymous):

But the answer is to be given in numbers

OpenStudy (anonymous):

Answer is -1

OpenStudy (vishweshshrimali5):

Yes and your answer is absolutely correct.

OpenStudy (anonymous):

How do I get that?

OpenStudy (vishweshshrimali5):

See: \[\tan(\cfrac{x_{2}}{2}) = \tan(\cfrac{x_{1} + \pi}{2})\] \[ = \tan(\cfrac{x_{1}}{2} + \cfrac{\pi}{2})\] \[= -\cfrac{1}{\tan(\cfrac{x_1}{2})}\] So, the required product will become -1.

OpenStudy (vishweshshrimali5):

Did you get it ?

OpenStudy (anonymous):

Yes thank you

OpenStudy (vishweshshrimali5):

No problem

OpenStudy (anonymous):

what is the value of tan pi/16 +2tan pi/8 +4

OpenStudy (anonymous):

can you'll help @vishweshshrimali5

OpenStudy (anonymous):

you*

OpenStudy (anonymous):

@amistre64 and @ganeshie8

OpenStudy (anonymous):

can youll help

OpenStudy (anonymous):

what is the value of tan pi/16 +2tan pi/8 +4

ganeshie8 (ganeshie8):

calculator not allowed ?

OpenStudy (anonymous):

no, the value is supposed to be in trgonometric ratios

OpenStudy (vishweshshrimali5):

Ok let me try

OpenStudy (vishweshshrimali5):

Ok try putting pi/16 as x

OpenStudy (vishweshshrimali5):

Also, 4 = 4 tan(pi/4)

OpenStudy (vishweshshrimali5):

So, expression will become, tan x + 2 tan 2x + 4 tan 4x

OpenStudy (vishweshshrimali5):

On simplifying this you will get, \[\cfrac{\sin(2x)}{\cos(2x) + 1} + \cfrac{2\sin(4x)}{\cos(4x) + 1} + \cfrac{4\sin(8x)}{\cos(8x) + 1}\] Put \(x = \cfrac{\pi}{16}\) You will get, \[\cfrac{\sin(2\cfrac{\pi}{16})}{\cos(2\cfrac{\pi}{16}) + 1} + \cfrac{2\sin(4\cfrac{\pi}{16})}{\cos(4\cfrac{\pi}{16}) + 1} + \cfrac{4\sin(8\cfrac{\pi}{16})}{\cos(8\cfrac{\pi}{16}) + 1}\] Solve and simplify it.

OpenStudy (vishweshshrimali5):

You will have to find out \(\sin(\cfrac{\pi}{8})\) and \(\cos(\cfrac{\pi}{8})\) which you can easily find out using the following formulae: (1) \(\sin{x} = \sqrt{\cfrac{1- \cos{2x}}{2}}\) (2) \(\cos{x} = \sqrt{\cfrac{1 + \cos{2x}}{2}}\)

OpenStudy (vishweshshrimali5):

I know its very tedious but that's all I can think at present. Sorry !

OpenStudy (anonymous):

olkay thankyou

OpenStudy (anonymous):

btw the answer has to come in trgo ratios... the answer is cot pi/16.... how ddo i get it though?